# Forecasts for Galaxy Formation and Dark Matter Constraints from Dwarf Galaxy Surveys

### v1

@@ 2.3 Modeling Subhalo Populations Beyond CDM
It is straightforward to extend the galaxy–halo connection model described above to certain beyond-CDM scenarios. In particular, our framework is based on halo properties measured at the time peak mass is achieved (for the luminosity model) or at infall (for the size and disruption models), and on halo orbits (for the disruption model). We do not expect any of these properties to change appreciably in models like WDM, which differ from CDM in terms of the initial conditions for structure formation, when studying halos well above the mass corresponding to the free-streaming scale, $M_{\mathrm{fs}}$. WDM constraints from observations of dwarf galaxies (e.g., Nadler et al. 2021c; Newton et al. 2021) and other small-scale structure probes (e.g., Banik et al. 2021b; Gilman et al. 2020a) imply that dwarf galaxies’ halos are orders-of-magnitude more massive than $M_{\mathrm{fs}}$, supporting this expectation. Furthermore, hydrodynamic WDM simulations indicate that dwarf galaxy formation is not significantly altered relative to CDM, even for models that are ruled out by current observations (e.g., Bozek et al. 2019; Fitts et al. 2019).
We therefore assume that our galaxy–halo connection model is unchanged in the beyond-CDM scenarios we consider, following Nadler et al. (2021c), and we focus on modeling variations in underlying halo abundances imprinted by new DM physics, as described in the following subsections.
@@ 2.3.1 Thermal-relic WDM
We consider thermal-relic WDM as a benchmark model with suppressed small-scale power and low-mass (sub)halo abundances (Bond & Szalay, 1983; Bode et al., 2001). The WDM particle mass $m_{\mathrm{WDM}}$, sets the free-streaming scale below which the linear matter power spectrum is suppressed. We parameterize this cutoff via (Nadler et al., 2021c)
@@
where the half-mode mass $M_{\mathrm{hm}}$ is directly related to the unknown WDM particle mass and corresponds (in linear theory) to the scale at which the linear matter power spectrum is suppressed by $75\%$ relative to CDM. [^4] Vogel & Abazajian (2023) find that Equation 4 is slightly inaccurate when compared to transfer functions output by linear Boltzmann solvers for $m_{\mathrm{WDM}}\gtrsim 3\ \mathrm{keV}$ (also see Appendix C of Decant et al. 2022). We do not correct for this discrepancy here because the (sub)halo mass function suppression we adopt is based on Equation 4; this will be addressed in a forthcoming study (E. O. Nadler et al., in preparation). We model the suppression of the SHMF using the fit to WDM zoom-in simulations from Lovell et al. (2014),
@@
where $f_{\mathrm{WDM}}\equiv(dn_{\mathrm{WDM}}/dM_{\mathrm{peak}})/(dn_{\mathrm{CDM}}/dM_{\mathrm{peak}})$ with $\alpha=2.7$, $\beta=1.0$, and $\gamma=-0.99$.
For the WDM analyses we present below, $M_{\mathrm{hm}}$ is a free parameter and each satellite is weighted by $f_{\mathrm{WDM}}\times(1-p_{\mathrm{disrupt}})$ when number counts are computed from realizations of our model. As discussed in Nadler et al. (2021c), our approach assumes that the WDM subhalo radial distribution is unchanged relative to CDM. WDM simulations indicate that this is a good assumption for subhalos more massive than $M_{\mathrm{hm}}$ (Lovell et al., 2021). Studying lower-mass subhalos would benefit from the direct use of WDM simulations, which we intend to pursue in future work.
@@ 2.3.2 Model-independent SHMF
where $f_{\mathrm{NCDM}}\equiv(dn_{\mathrm{NCDM}}/dM_{\mathrm{peak}})/(dn_{\mathrm{CDM}}/dM_{\mathrm{peak}})$.
Thus, $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ parameterize SHMF deviations from CDM centered on peak masses of $10^{8}$, $10^{9}$, and $10^{10}~{}M_{\mathrm{\odot}}$, respectively. For our SHMF forecasts, $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ are free parameters and each satellite is weighted by $f_{\mathrm{NCDM}}\times(1-p_{\mathrm{disrupt}})$ when number counts are computed from realizations of our model. Like our WDM model, our prescription for modifying the SHMF does not alter other aspects of subhalo populations (e.g., radial distributions) and does not couple to the rest of our galaxy–halo connection parameters. Although these assumptions may be broken in specific beyond-CDM scenarios, this approach allows us to disentangle the effects of variations in the SHMF and galaxy–halo connection.
@@ 2.4 Model Illustration
Before performing these full parameter recovery tests, we illustrate the predictions of our model and the constraining power of satellite population measurements. Figure 1 shows mock observations of mean satellite luminosity functions from our forty MW-mass zoom-in simulations. We compare mock observations of “true” satellite luminosity functions in a model with $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$ (shown as error bars) to predictions with $\mathcal{M}_{50}=3\times 10^{7}$, $10^{8}$, and $3\times 10^{8}~{}M_{\mathrm{\odot}}$ (shown as lines), assuming that all satellites of each host are detectable. Below, we refer to $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$ ($3\times 10^{7}~{}M_{\mathrm{\odot}}$) as a “strong” (“weak”) cutoff scenario, and we set $\sigma_{M}=0.2~{}\mathrm{dex}$, and $\mathcal{S}_{\mathrm{gal}}=0.2$ in both scenarios; see Section 4 for details. Note that all remaining galaxy–halo connection parameters are fixed to the input values listed above.
The model with $\mathcal{M}_{50}=3\times 10^{8}~{}M_{\mathrm{\odot}}$ is detectable using the complete satellite population of a single MW-mass host. Meanwhile, models with $\mathcal{M}_{50}\lesssim 10^{8}~{}M_{\mathrm{\odot}}$ are more difficult to detect because they only affect the abundances of the faintest satellites we consider. For comparison, the gray points in Figure 1 show the luminosity function of kinematically confirmed and candidate MW satellites detected by DES and PS1 (Drlica-Wagner et al., 2020). Because MW satellite observations are incomplete, we also show the completeness-corrected estimate from Drlica-Wagner et al. (2020) for comparison. Note that Nadler et al. (2020b) infer $\mathcal{M}_{50}<8.5\times 10^{7}~{}M_{\mathrm{\odot}}$ at $95\%$ confidence using these satellites, including the observational selection function derived in Drlica-Wagner et al. (2020).
Figure 1 also shows a “weak” galaxy formation cutoff combined with a $m_{\mathrm{WDM}}=6.5~{}\mathrm{keV}$ WDM model, corresponding to the $95\%$ confidence WDM limit from the MW satellite population derived in Nadler et al. (2021c). These astrophysical and DM models combine to yield a luminosity function similar to a “strong” galaxy formation cutoff in CDM, illustrating how degeneracies between galaxy formation and DM physics can manifest in dwarf galaxy populations.
@@
Figure 1: Mean satellite luminosity function and $1\sigma$ Poisson uncertainty for the MW satellite population observed by DES and PS1 (gray), and a completeness-corrected total luminosity function from (Drlica-Wagner et al. 2020; black). Blue error bars show the predicted luminosity function mean and $1\sigma$ Poisson uncertainty for forty complete satellite populations of MW-mass hosts generated from our galaxy–halo connection model with a “strong” galaxy formation cutoff ($\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$, $\sigma_{M}=0.2~{}\mathrm{dex}$, and $\mathcal{S}_{\mathrm{gal}}=0.2$; all remaining galaxy–halo connection parameters are fixed to the input values listed in Section 2.4). Blue lines show predictions for the same $\sigma_{M}$ and $\mathcal{S}_{\mathrm{gal}}$, with $\mathcal{M}_{50}=3\times 10^{8}$ (dashed), $10^{8}$ (solid), and $3\times 10^{7}~{}M_{\mathrm{\odot}}$ (dot-dashed, which we refer to as a “weak” cutoff throughout). The blue band shows $16$th–$84$th percentile host-to-host scatter for $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$. All predictions assume a satellite detectability threshold of $M_{V}<0~{}\mathrm{mag}$ and $\mu<32~{}\mathrm{mag\ arcsec}^{-2}$, and error bars are offset horizontally for clarity.
@@
Figure 2: Left: suppression of the subhalo mass function (defined as the ratio of the WDM to CDM SHMF) for WDM models with $m_{\mathrm{WDM}}=4.9$, $5.6$, $7.9$, and $17.0~{}\mathrm{keV}$ from lightest to darkest red. Our “strong” and “weak” galaxy formation cutoff scenarios are shown on the same axes. The blue band shows the range of $\mathcal{M}_{50}$ probed at $68\%$ confidence by our “strong” galaxy formation cutoff forecast, and thus illustrates the range of halo masses we typically probe. Right: mean SHMF (black) from our forty CDM MW-mass zoom-in simulations, measured using peak virial mass. Dashed blue, dot-dashed purple, and dotted red lines illustrate the effects of our model-independent SHMF parameterization, which can enhance or suppress the SHMF relative to CDM as a function of peak halo mass, depending on the values of $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$.
@@
Finally, the right panel of Figure 2 illustrates our model-independent SHMF parameterization. We show suppressed (red) and enhanced (blue) models constrained by one complete satellite population in our forecasts below (see Section 6 for details), along with a model that mixes SHMF suppression and enhancement in different peak virial mass decades (purple). Our model can alter both the shape and normalization of the SHMF; only some of these deviations can be mimicked by galaxy formation or WDM cutoffs.
@@ 3.1 Observational Assumptions
The second assumption is implicit in our procedure and follows because we compare predicted satellite populations, generated using the subhalo population around a given simulated host, to a “true” satellite population generated from the same host. Both assumptions are optimistic; we make these choices in order to assess the potential of dwarf satellite populations for constraining galaxy formation and DM physics in an optimal scenario, and we discuss caveats in Section 7.
@@ 3.2 Likelihood
To fit the mock data and assess parameter uncertainties, we generalize the likelihood framework from Nadler et al. (2019, 2020b), following Nadler (2021). In particular, we model realizations of satellite luminosities and sizes as a Poisson process, drawn from a distribution uniquely predicted by a set of parameter values $\bm{\theta}=\{\alpha,\sigma_{M},\mathcal{A},n,\sigma_{\log R},\mathcal{M}_{50},\mathcal{S}_{\mathrm{gal}},\mathcal{B}\}$, and additionally $M_{\mathrm{hm}}$ or $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ in our beyond-CDM forecasts. The probability of detecting $n_{\mathrm{obs},ij}$ satellites in luminosity and size bin $j$, in host $i$, is given by
@@
where ${N_{\mathrm{real.}}}$ is the total number of realizations per host at a given $\bm{\theta}$, ${N}_{ij}\equiv\sum_{k=1}^{{N_{\mathrm{real.}}}}n_{\mathrm{pred},ijk}$, and $n_{\mathrm{pred},ijk}$ is the number of predicted satellites in host $i$, bin $j$, and realization $k$. Equation 7 is a modified Poisson likelihood (Nadler et al., 2019)
@@
derived by marginalizing over the (unknown) underlying Poisson rate in each bin of observable parameter space. In practice, we replace the factorials in Equation 8 with appropriate Gamma functions because our model produces non-integer satellite counts.
Evaluations of Equations 7 and 8 include $N_{\mathrm{real.}}$ realizations at a given $\bm{\theta}$ because our galaxy–halo connection model is stochastic. In this study, we use $N_{\mathrm{real.}}=10$ per host at a given $\bm{\theta}$. We have checked that varying $N_{\mathrm{real.}}$ does not impact our results; in particular, our Markov Chain Monte Carlo (MCMC) evaluation of the posterior converges in fewer steps when $N_{\mathrm{real.}}$ is larger, but the posterior distribution and total run time are unchanged. We follow Nadler et al. (2020b) by using $14$ bins in $M_{V}$, evenly spaced from 00 to $-20\ \mathrm{mag}$, and two bins in surface brightness that split the sample at $\mu_{V}=28\ \mathrm{mag\ arcsec}^{-2}$.
@@
3.3 Posterior Sampling
It is difficult to sample our full galaxy–halo connection (plus beyond-CDM) parameter space using more than two hosts at a time. Because our model is evaluated on high-resolution cosmological simulations, likelihood evaluations are computationally expensive when using large numbers of hosts. In addition, degeneracies among galaxy–halo connection parameters (and, in our WDM and SHMF forecasts, between galaxy–halo connection and beyond-CDM parameters) slow down posterior sampling. The most challenging parameter in this context is $\sigma_{M}$, which is often degenerate with $\mathcal{M}_{50}$, $\mathcal{S}_{\mathrm{gal}}$, and beyond-CDM parameters that mimic a cutoff; we discuss specific degeneracies in each forecast below. Furthermore, because the SHMF rises steeply with decreasing subhalo mass, abundant low-mass subhalos preferentially up-scatter to observable luminosities, resulting in non-Gaussian posteriors for individual hosts. The complexity of the fit increases with the number of hosts used to compute the joint likelihood since each host has a different underlying subhalo population and associated non-Gaussian posterior.

### v2

@@ 2.3 Modeling Subhalo Populations Beyond CDM
The galaxy–halo connection model described in Section 2.2 can easily be applied to beyond-CDM scenarios. Here, we rely on subhalo populations from our CDM simulations (Section 2.1) to analyze WDM and models with a generalized SHMF. Thus, our analyses of these scenarios rely on the assumption that the following input subhalo properties do not significantly differ relative to CDM: $V_{\mathrm{peak}}$ (for the luminosity model), $R_{\mathrm{vir,acc}}$ (for the size model), $M_{\mathrm{peak}}$ (for the occupation fraction model), and (mainly) accretion time and orbital properties (for the disruption model). In addition, the radial and angular distribution of subhalos is taken directly from our CDM simulations. We justify these assumptions for the WDM and generalized SHMF cases below.
@@ 2.3.1 Thermal-relic WDM
We consider thermal-relic WDM as a benchmark model with suppressed small-scale power and low-mass (sub)halo abundances (Bond & Szalay, 1983; Bode et al., 2001). The WDM particle mass, $m_{\mathrm{WDM}}$, sets the free-streaming scale below which the linear matter power spectrum is suppressed. We parameterize this cutoff via (Nadler et al., 2021c)
@@
where the half-mode mass, $M_{\mathrm{hm}}$, is related to the unknown WDM particle mass, $m_{\mathrm{WDM}}$. In particular, $M_{\mathrm{hm}}$ is defined by the wavenumber at which the linear matter power spectrum is suppressed by $75\%$ relative to CDM (Nadler et al., 2019a). Thus, theoretical uncertainties in Equation 4 are small, and we hold this relation fixed in our forecasts. [^4] Vogel & Abazajian (2023) found that Equation 4 is slightly inaccurate, at the $\approx 10\%$ level, when compared to transfer functions output by linear Boltzmann solvers for $m_{\mathrm{WDM}}\gtrsim 3\ \mathrm{keV}$ (also see Appendix C of Decant et al. 2022). We do not correct for this discrepancy here because the (sub)halo mass function suppression we adopt is based on Equation 4; this will be addressed in a forthcoming study (E. O. Nadler et al. 2024, in preparation).
We model the suppression of the SHMF using the fit to WDM zoom-in simulations from Lovell et al. (2014),
@@
where $f_{\mathrm{WDM}}\equiv(dn_{\mathrm{WDM}}/dM_{\mathrm{peak}})/(dn_{\mathrm{CDM}}/dM_{\mathrm{peak}})$ with $\alpha=2.7$, $\beta=1.0$, and $\gamma=-0.99$. For the WDM analyses we present below, $M_{\mathrm{hm}}$ is a free parameter and each satellite is weighted by $f_{\mathrm{WDM}}\times(1-p_{\mathrm{disrupt}})$ when number counts are computed from realizations of our model.
We focus on models with $m_{\mathrm{WDM}}\gtrsim 5~{}\mathrm{keV}$ ($M_{\mathrm{hm}}\gtrsim 10^{8}~{}M_{\mathrm{\odot}}$). As a result, we primarily study halos above the half-mode mass (and well above the free-streaming mass) in our forecasts. In this regime, WDM simulations indicate that $M_{\mathrm{peak}}$ and $V_{\mathrm{peak}}$ are only reduced by $\mathcal{O}(10\%)$ relative to matched CDM halos (e.g., Bozek et al. 2019; Fitts et al. 2019). We expect $R_{\mathrm{vir,acc}}$ to be reduced by a similar amount because it is mainly determined by $M_{\mathrm{peak}}$ and the amount of pre-infall stripping a subhalo experiences, which does not differ significantly between the models (e.g., Elahi et al. 2014). Meanwhile, as described in Section 2.2, orbital properties of given WDM subhalos are not significantly altered relative to CDM.
Finally, we assume that the WDM subhalo radial distribution is unchanged relative to CDM. WDM simulations of MW-mass hosts indicate that this is a good assumption for subhalos more massive than $M_{\mathrm{hm}}$ (Lovell et al., 2021), which we primarily study. Modeling lower-mass subhalos would benefit from the direct use of WDM simulations. Thus, our method for applying the model to WDM is well motivated, and we leave an extension of our framework to beyond-CDM simulations for future work.
@@ 2.3.2 Model-independent SHMF
where $f_{\mathrm{NCDM}}\equiv(dn_{\mathrm{NCDM}}/dM_{\mathrm{peak}})/(dn_{\mathrm{CDM}}/dM_{\mathrm{peak}})$. In this model, $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ parameterize SHMF deviations from CDM centered on peak masses of $10^{8}$, $10^{9}$, and $10^{10}~{}M_{\mathrm{\odot}}$, respectively. For our SHMF forecasts, $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ are free parameters, and each satellite is weighted by $f_{\mathrm{NCDM}}\times(1-p_{\mathrm{disrupt}})$ when number counts are computed from realizations of our model.
By construction, our generalized SHMF scenarios do not alter any of the subhalo properties relevant for our galaxy–halo connection model or quantities like the radial distribution. Although these assumptions may be broken in specific beyond-CDM scenarios, this approach allows us to isolate the effects of varying the SHMF.
@@ 2.4 Model Illustration
Figure 1: Mean satellite luminosity function and $1\sigma$ Poisson uncertainty for the MW satellite population observed by DES and PS1 (gray), and a completeness-corrected total luminosity function from (Drlica-Wagner et al. 2020; black). Blue error bars show the predicted luminosity function mean and $1\sigma$ Poisson uncertainty for $40$ complete satellite populations of MW-mass hosts generated from our galaxy–halo connection model with a “strong” galaxy formation cutoff ($\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$, $\sigma_{M}=0.2~{}\mathrm{dex}$, and $\mathcal{S}_{\mathrm{gal}}=0.2$; all remaining galaxy–halo connection parameters are fixed to the input values listed in Section 2.4). Blue lines show predictions for the same $\sigma_{M}$ and $\mathcal{S}_{\mathrm{gal}}$, with $\mathcal{M}_{50}=3\times 10^{8}$ (dashed), $10^{8}$ (solid), and $3\times 10^{7}~{}M_{\mathrm{\odot}}$ (dotted–dashed, which we refer to as a “weak” cutoff throughout). The blue band shows $16$th–$84$th percentile host-to-host scatter for $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$. All predictions assume a satellite detectability threshold of $M_{V}<0~{}\mathrm{mag}$ and $\mu<32~{}\mathrm{mag\ arcsec}^{-2}$, and error bars are offset horizontally for clarity.
@@
Figure 2: Left: suppression of the subhalo mass function (defined as the ratio of the WDM to CDM SHMF) for WDM models with $m_{\mathrm{WDM}}=4.9$, $5.6$, $7.9$, and $17.0~{}\mathrm{keV}$ from lightest to darkest red. Our “strong” and “weak” galaxy formation cutoff scenarios are shown on the same axes. The blue band shows the range of $\mathcal{M}_{50}$ probed at $68\%$ confidence by our “strong” galaxy formation cutoff forecast, and thus illustrates the range of halo masses we typically probe. Right: mean SHMF (black) from our $40$ CDM MW-mass zoom-in simulations, measured using peak virial mass. Dashed blue, dotted–dashed purple, and dotted red lines illustrate the effects of our model-independent SHMF parameterization, which can enhance or suppress the SHMF relative to CDM as a function of peak halo mass, depending on the values of $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$. The uniformly suppressed (dashed blue) and enhanced (dotted red) models correspond to our projections for $68\%$ confidence SHMF constraints from one complete host.
Before performing these full parameter recovery tests, we illustrate the predictions of our model and the constraining power of satellite population measurements. Figure 1 shows mock observations of mean satellite luminosity functions from our $40$ MW-mass zoom-in simulations. We compare mock observations of “true” satellite luminosity functions in a model with $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$ (shown as error bars) to predictions with $\mathcal{M}_{50}=3\times 10^{7}$, $10^{8}$, and $3\times 10^{8}~{}M_{\mathrm{\odot}}$ (shown as lines). Below, we refer to $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$ ($3\times 10^{7}~{}M_{\mathrm{\odot}}$) as a “strong” (“weak”) cutoff scenario, and we set $\sigma_{M}=0.2~{}\mathrm{dex}$, and $\mathcal{S}_{\mathrm{gal}}=0.2$ in both scenarios; see Section 4 for details. Note that all remaining galaxy–halo connection parameters are fixed to the input values listed above.
For this illustration, we optimistically assume that all satellites of each host with $M_{V}<0~{}\mathrm{mag}$ and $\mu_{V}<32~{}\mathrm{mag\ arcsec}^{-2}$ are detectable; we expand on our observational assumptions in Section 3.1 and discuss them in the context of near and long-term observational capabilities in Sections 7.1 and 7.2. Given this assumption, the model with $\mathcal{M}_{50}=3\times 10^{8}~{}M_{\mathrm{\odot}}$ is detectable using the complete satellite population of a single MW-mass host. Meanwhile, models with $\mathcal{M}_{50}\lesssim 10^{8}~{}M_{\mathrm{\odot}}$ are more difficult to detect because they only affect the abundances of the faintest satellites we consider. For comparison, the gray points in Figure 1 show the luminosity function of kinematically confirmed and candidate MW satellites detected by DES and PS1 (Drlica-Wagner et al., 2020). Because MW satellite observations are incomplete, we also show the completeness-corrected estimate from Drlica-Wagner et al. (2020) for comparison. Note that Nadler et al. (2020b) inferred $\mathcal{M}_{50}<8.5\times 10^{7}~{}M_{\mathrm{\odot}}$ at $95\%$ confidence using these satellites, including the observational selection function derived in Drlica-Wagner et al. (2020).
Figure 1 also shows a “weak” galaxy formation cutoff combined with an $m_{\mathrm{WDM}}=6.5~{}\mathrm{keV}$ WDM model, corresponding to the $95\%$ confidence WDM limit from the MW satellite population derived in Nadler et al. (2021c). These astrophysical and DM models combine to yield a luminosity function similar to a “strong” galaxy formation cutoff in CDM, illustrating how degeneracies between galaxy formation and DM physics can manifest in dwarf galaxy populations.
@@
Finally, the right panel of Figure 2 illustrates our model-independent SHMF parameterization. We show uniformly suppressed (dotted red) and enhanced (dashed blue) models constrained at $68\%$ confidence by one complete satellite population in our forecasts; we also show a model that mixes SHMF suppression and enhancement in different peak virial mass decades (dotted–dashed purple) at a level that is consistent with our projected dwarf galaxy sensitivity (see Section 6.1 for details). For context, the size of these SHMF deviations relative to CDM is comparable to the SHMF sensitivity inferred from recent stellar stream and strong lensing measurements (Gilman et al., 2020a; Banik et al., 2021a). Both the shape and normalization of the SHMF can be altered in our parameterization, and only some of this parameter space can be mimicked by galaxy formation or WDM cutoffs.
@@ 3.1 Observational Assumptions
The second assumption is implicit in our procedure and follows because we compare predicted satellite populations, generated using the subhalo population around a given simulated host, to a “true” satellite population generated from the same host. Both assumptions are optimistic; we make these choices in order to assess the potential of dwarf satellite populations for constraining galaxy formation and DM physics. We place these assumptions in the context of current and future observational sensitivity and discuss related areas for modeling work in Section 7.
@@ 3.2 Likelihood
To fit the mock data and assess parameter uncertainties, we generalize the likelihood framework from Nadler et al. (2019b, 2020b), following Nadler (2021). In particular, we model realizations of satellite luminosities and sizes as a Poisson process, drawn from a distribution uniquely predicted by a set of parameter values $\bm{\theta}=\{\alpha,\sigma_{M},\mathcal{A},n,\sigma_{\log R},\mathcal{M}_{50},\mathcal{S}_{\mathrm{gal}},\mathcal{B}\}$, and additionally $M_{\mathrm{hm}}$ or $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ in our beyond-CDM forecasts. The probability of detecting $n_{\mathrm{obs},ij}$ satellites in luminosity and size bin $j$, in host $i$, is given by
@@
where ${N_{\mathrm{real.}}}$ is the total number of realizations per host at a given $\bm{\theta}$, ${N}_{ij}\equiv\sum_{k=1}^{{N_{\mathrm{real.}}}}n_{\mathrm{pred},ijk}$, and $n_{\mathrm{pred},ijk}$ is the number of predicted satellites in host $i$, bin $j$, and realization $k$. In practice, we replace the factorials in Equation 7 with appropriate Gamma functions because our model produces noninteger satellite counts.
Before moving on, we provide context for this likelihood. In particular, Equation 7 is a modified Poisson likelihood derived by marginalizing over the (unknown) underlying Poisson rate in each bin of observable parameter space via (Nadler et al., 2019b)
@@
Here, the final expression is obtained by substituting Poisson likelihoods for each conditional probability and a flat prior on positive values of $\lambda_{ij}$. The key assumption is that mock and observed satellites are drawn from the same (unknown) Poisson distribution; our likelihood converges to this distribution in the limit of many realizations (Nadler et al., 2019b).
When evaluating Equation 7, we include $N_{\mathrm{real.}}$ realizations at a given $\bm{\theta}$ because our galaxy–halo connection model is stochastic. In this study, we use $N_{\mathrm{real.}}=10$ per host at a given $\bm{\theta}$. We have checked that varying $N_{\mathrm{real.}}$ does not impact our results; in particular, our Markov Chain Monte Carlo (MCMC) evaluation of the posterior converges in fewer steps when $N_{\mathrm{real.}}$ is larger, but the posterior distribution and total run time are unchanged. We follow Nadler et al. (2020b) by using $14$ bins in $M_{V}$, evenly spaced from 00 to $-20\ \mathrm{mag}$, and two bins in surface brightness that split the sample at $\mu_{V}=28\ \mathrm{mag\ arcsec}^{-2}$.
@@
3.3 Posterior Sampling
It is difficult to sample our full galaxy–halo connection (plus beyond-CDM) parameter space using more than two hosts at a time. Because our model is evaluated on high-resolution cosmological simulations, likelihood evaluations are computationally expensive when using large numbers of hosts. In addition, degeneracies among galaxy–halo connection parameters (and, in our WDM and SHMF forecasts, between galaxy–halo connection and beyond-CDM parameters) slow down posterior sampling. The most challenging parameter in this context is $\sigma_{M}$, which is often degenerate with $\mathcal{M}_{50}$, $\mathcal{S}_{\mathrm{gal}}$, and beyond-CDM parameters that mimic a cutoff; we discuss specific degeneracies in each forecast below. Furthermore, because the SHMF rises steeply with decreasing subhalo mass, abundant low-mass subhalos preferentially up-scatter to observable luminosities (a form of Eddington bias), resulting in non-Gaussian posteriors for individual hosts. The complexity of the fit increases with the number of hosts used to compute the joint likelihood since each host has a different underlying subhalo population and associated non-Gaussian posterior.